Find a linear homogeneous differential equation, whose solutions are {e^x,e^-x,e^x-2e^-x}
1
Expert's answer
2020-12-20T18:41:33-0500

Let us solve the linear homogeneous differential equation "y''-y=0". The characteristic equation "k^2-1=0" has the solutions "k_1=1" and "k_2=-1". Therefore, the general solution is "y=C_1e^x+C_2e^{-x}". If "C_1=1" and "C_2=0" we have "y=e^x". If "C_1=0" and "C_2=1" we have we have "y=e^{-x}". If "C_1=1" and "C_2=-2" we have "y=e^x-2e^{-x}."


Answer: "y''-y=0"